# Dolphin: A Programmable Framework for Scalable Neurosymbolic Learning

### v3

@@ Abstract
Neurosymbolic learning enables the integration of symbolic reasoning with deep learning but faces significant challenges in scaling to complex symbolic programs, large datasets, or both. We introduce Dolphin, a framework that tackles these challenges by supporting neurosymbolic programs in Python, executing complex symbolic reasoning on the CPU while vectorizing probabilistic computations and gradient propagation on the GPU. Across 13 benchmarks spanning tasks over text, image, and video data, with symbolic reasoning features like recursion and black-box functions, Dolphin converges to state-of-the-art accuracies on the more complex benchmarks while existing frameworks such as Scallop, ISED, and IndeCateR+ fail to converge within the time limit. On simpler benchmarks, Dolphin matches their performance, while achieving these results 1.71x to 62x faster than the baselines. Overall, Dolphin advances the scalability of neurosymbolic frameworks, achieving state-of-the-art efficiency and convergence on difficult benchmarks where existing frameworks struggle.
@@ 1 Introduction
In this paper, we propose Dolphin as a solution for scaling neurosymbolic learning. In Dolphin, we build three key components that effectively tackle scalability challenges with existing neurosymbolic frameworks. First, we develop a general symbolic representation that efficiently captures the relationships between neural network outputs as PyTorch tensors on GPU and associated discrete symbols as Python objects on CPU. Second, we introduce a set of primitives to enable writing symbolic manipulations that can be mapped to computations over these representations, while allowing support for black-box Python functions that simplify the writing of complex symbolic programs. Third, we develop a set of vectorized provenance semirings (Green et al., 2007) that are easily pluggable into Dolphin and enable to efficiently compute symbolic gradients.
Together, these components allow Dolphin to build a computation graph that spans both the neural and probabilistic computations () that is highly parallelizable, and end-to-end differentiable on GPU. At the same time, it runs over discrete symbols on CPU to better manage GPU memory. This allows Dolphin to scale effectively to complex problems such as CLUTRR-N (Table 2) where denotes the max length of the reasoning chain in the training dataset. In the case of Scallop, as the length of the reasoning chain increases, the gap between convergence times dramatically widens. In contrast, for , Dolphin takes around 13 minutes to converge, about 5x faster than Scallop, while for , Dolphin takes around 15 minutes, about 8.5x faster than Scallop. Finally, Dolphin is implemented as a library integrated with PyTorch, allowing users to easily incorporate it into their existing deep learning pipelines.
@@ 2.1.2 Operations
Since the tags here are PyTorch tensors representing probabilities, the implementations of the conjunction and disjunction operations are dictated by the underlying provenance specified by the program. A more detailed explanation of the provenances is provided in Section 2.2.
@@ 3.2 Experimental Setup and Baselines
Setup. All experiments, except CLUTRR, were run on machines with two 20-core Intel Xeon Gold 6248 CPUs, four NVIDIA GeForce RTX 2080 Ti (11 GB) GPUs, and 768 GB RAM. Since CLUTRR demands more GPU memory due to running the RoBERTa model with a standard batch size of 16, all tools were run with a NVIDIA A100 40GB GPU. We ran each tool thrice until convergence or until a timeout of 10 hours was reached and report the average best accuracy and training time.
@@ 3.3 RQ1: Scalability
Table 2 presents the total training times () till convergence in seconds for Dolphin and Scallop across all benchmarks, alongside the scaling factor (the ratio of the total training times of the baselines to Dolphin). Table 3 shows the same for the remaining baselines over SumN and HWF. The results demonstrate that Dolphin advances the state-of-the-art in neurosymbolic learning by scaling to more complex problems, e.g., larger versions of HWF, Path, and Mugen, that are beyond the reach of the other baselines which time out within 10 hours. Even for the other benchmarks where baselines do not time out, Dolphin achieves a scalability factor of up to 47x, 62x, and 3.49x against Scallop, ISED, and IndeCateR+ with an average speed up of 13.95x across all baselines for all benchmarks.
When the baselines do timeout, we use the training times per epoch to calculate a scalability factor (Table 5, Appendix D). We see that among cases where baselines timeout, is up to 280x for HWF-19 and is 36.4x faster on average. This results in Dolphin effectively training for more epochs in less time compared with the other baselines, which also allows it to converge to higher accuracies, as we see in Figure 3. We expand on these results in the next RQ.
@@ 3.4 RQ2: Accuracy
Figure 3 presents the accuracy of Dolphin and the baselines on the different benchmarks. Dolphin accuracies are marked in blue. In all cases, for Dolphin, we report the accuracies of the best-performing provenance. We use the DAMP provenance for MNIST, CLUTRR, and Mugen benchmarks, and the DTKP-AM provenance for the HWF and PathFinder benchmarks.
We observe that in all cases, Dolphin achieves state-of-the-art accuracy among general-purpose neurosymbolic frameworks, except in CLUTRR-3, where Dolphin’s accuracy is slightly lower than Scallop’s. As we scale up to larger versions of the benchmarks, Dolphin achieves better accuracy, because the baselines either report lower accuracy due to the complexity of the benchmark (e.g., black-box sampling techniques such as ISED/IndeCateR+ on HWF) or fail to converge within 10 hours due to slower per-epoch train time (e.g., Scallop on PathFinder-256). As a result, Dolphin significantly outperforms the second-best benchmark on the largest dataset versions, achieving a 20% gain on HWF, 33% on Path, and up to 38% on Mugen. Most importantly, these results show that not only do Dolphin’s scalability improvements not come at the cost of accuracy, but Dolphin enables SOTA accuracy when previously unattainable.
@@ 4 Related Work
Scaling techniques. Various methods exist to scale differentiable reasoning. LYRICS (Marra et al., 2019), Logic Tensor Networks (Badreddine et al., 2022), and Tensorlog (Cohen et al., 2020) compile first-order logic constraints into GPU-compatible computation graphs. Other techniques, such as Greedy NTP (Minervini et al., 2020a) and the conditional theorem prover (Minervini et al., 2020b), optimize proof search using heuristics. However, these methods are often task-specific and lack generalizability to broader neurosymbolic learning, or they struggle to scale due to memory constraints when grounding symbolic computations on GPU.
Specialized neurosymbolic solutions. There are many specialized solutions for various neurosymbolic tasks. For instance, NGS (Li et al., 2020) uses a hand-coded syntax to specify the structure of mathematical expressions for HWF. More general solutions, such as NS-CL (Mao et al., 2019) includes a framework for visual question answering that learns symbolic representations for text and images. NeRd (Chen et al., 2021) transforms questions in natural language into executable programs based on symbolic information extracted from text. (Orvieto et al., 2023) proposes a recurrent neural network architecture that achieves 95% accuracy on Path (S) and 94% on Path (M). In contrast, Dolphin is a general system that tries to scale diverse neurosymbolic programs.
@@ 5 Conclusion and Limitations
We proposed Dolphin, a system for scaling neurosymbolic learning. Dolphin provides abstractions for writing symbolic programs along with pluggable vectorized provenances to compute symbolic gradients. This allows users to write differentiable symbolic programs in Python within PyTorch pipelines that can scale to complex programs and large datasets. We show that Dolphin scales significantly better than existing neurosymbolic frameworks while achieving state-of-the-art performance on a variety of tasks.
@@ A.1 WMC approximation
Under both and the add-mult approximation, the probability of is:
@@ D.1 Comparison of Per Epoch Training Times
Dolphin trains slightly slower than Scallop on both versions of the Mugen benchmark. This is because the Dolphin program written for Mugen uses Python objects and operations that are not fully batchable across samples. In contrast, the Scallop program, which is written in a compiled and optimized language, runs around 1.3x faster than Dolphin on average per iteration. However, Dolphin requires only 20 epochs to converge whereas Scallop requires almost 1000 epochs (Li et al., 2023). The accuracy curves are shown and discussed in 6 and 7. As a result, Dolphin’s total training time is still lower than Scallop’s (3.5x for Mugen-1K).
D.2 MNIST Sum-N
For this task, the base neural network model is a standard CNN classifying each image into 10 classes of digits (0, 1, , 9). The symbolic module sums the Distribution objects over the logits output by the neural model for each image.
@@ D.2 MNIST Sum-N
D.3 Hand-Written Formula
@@ D.6 Mugen
D.7 Accuracy Curves for Mugen
Figures 6 and 7 show the test accuracies for each epoch of Dolphin and Scallop for Mugen (Small) and Mugen (Medium) respectively. The accuracies for TVR are on the left and VTR are on the right. Note that Dolphin converges within 20 epochs in all cases, while Scallop does not converge within 100 epochs. Since Dolphin and Scallop implement identical symbolic programs and neural network components, we attribute Dolphin’s better convergence to differences in the neurosymbolic backend. Importantly, the two use different frameworks for differentiable symbolic reasoning: Dolphin uses PyTorch while Scallop uses its own implementation of auto-differentiation. There could be optimizations in PyTorch that are absent from Scallop, where the performance boost is magnified by the relative complexity of Mugen compared to the other tasks. However, this claim requires further investigation to confirm.

### v4

@@ Abstract
Neurosymbolic learning enables the integration of symbolic reasoning with deep learning but faces significant challenges in scaling to complex symbolic programs, large datasets, or both. We introduce Dolphin, a framework that tackles these challenges by supporting neurosymbolic programs in Python, executing complex symbolic reasoning on the CPU while vectorizing probabilistic computations and gradient propagation on the GPU. Across 13 benchmarks spanning tasks over text, image, and video data, with symbolic reasoning features like recursion and black-box functions, Dolphin converges to state-of-the-art accuracies on the more complex benchmarks while existing frameworks such as Scallop, ISED, and IndeCateR+ fail to converge within the time limit. On simpler benchmarks, Dolphin matches their performance, while achieving these results 1.71x to 62x faster than the baselines. Overall, Dolphin advances the scalability of neurosymbolic frameworks, achieving state-of-the-art efficiency and convergence on difficult benchmarks where existing frameworks struggle. The code is published at https://github.com/Dolphin-NeSy/Dolphin.
@@ 1 Introduction
In this paper, we propose Dolphin as a solution for scaling neurosymbolic learning. In Dolphin, we build three key components that effectively tackle scalability challenges with existing neurosymbolic frameworks. First, we develop a unified representation that efficiently captures the relationships between neural network outputs as PyTorch tensors on GPU and associated discrete symbols as Python objects on CPU. Second, we introduce a set of primitives to enable writing symbolic manipulations that can be mapped to computations over these representations, while allowing support for black-box Python functions that simplify the writing of complex symbolic programs. Third, we develop a set of vectorized provenance semirings (Green et al., 2007) that are easily pluggable into Dolphin and enable to efficiently compute symbolic gradients.
Together, these components enable Dolphin to construct a computation graph that integrates both neural and probabilistic computations (), ensuring high parallelism and end-to-end differentiability on GPU. At the same time, it runs over discrete symbols on CPU, allowing flexible manipulation over arbitrary Python objects. This allows Dolphin to scale effectively to complex problems such as CLUTRR-N (Table 2) where denotes the max length of the reasoning chain in the training dataset. In the case of Scallop, as the length of the reasoning chain increases, the gap between convergence times dramatically widens. In contrast, for , Dolphin takes around 13 minutes to converge, about 5x faster than Scallop, while for , Dolphin takes around 15 minutes, about 8.5x faster than Scallop. Finally, Dolphin is implemented as a library integrated with PyTorch, allowing users to easily incorporate it into their existing deep learning pipelines.
@@ 2.1.1 Abstractions
Example 2.1.
Consider MNIST images for the SumN task discussed in Section LABEL:sec:overview. Let be the neural model which classifies image into one of 10 classes representing digits 0 to 9. Let and . We thus define the following Distributions:
1.
2.
@@ 2.1.2 Operations
Since the tags here are PyTorch tensors representing probabilities, the implementations of the conjunction and disjunction operations are dictated by the underlying provenance specified by the program, detailed in Section 2.2.
Example 2.2.
Continuing from Example 2.1, let function be applied to and to produce a new Distribution . thus represents a Distribution over the sum of the symbols from and :
Consider the tag of . 1 can be a result of being 0 and being 1, or of being 1 and being 0:
This expression’s value (e.g. 0.70) depends on the provenance specified (e.g. DAMP), discussed in Section 2.2.
@@
Example 2.3.
Continuing from Example 2.1, assume we want a Distribution over just the even symbols of . We can consider a filtering function . The resulting Distribution will have all the odd-numbered symbols completely removed:
@@ 3.2 Experimental Setup and Baselines
Setup. All experiments, except CLUTRR, were run on machines with two 20-core Intel Xeon Gold 6248 CPUs, four NVIDIA GeForce RTX 2080 Ti (11 GB) GPUs, and 768 GB RAM. Since CLUTRR demands more GPU memory due to running the RoBERTa model with a standard batch size of 16, all programs for this benchmark were run with a NVIDIA A100 40GB GPU. We ran each tool thrice until convergence or until a soft timeout of 10 hours was reached and report the average best accuracy and training time. For HWF/MNIST, we use the same CNN architecture as Scallop (Appendix D). For CLUTRR, we use Scallop’s Roberta configuration: a pretrained model (roberta-base) finetuned while training the classification head.
@@ 3.3 RQ1: Scalability
Table 2 presents the total training times () till convergence in seconds for Dolphin and Scallop across all benchmarks, alongside the scaling factor (the ratio of the total training times of the baselines to Dolphin). Table 3 shows the same for the remaining baselines over SumN and HWF. We set a soft timeout of 10 hours, though we still report the training times that run over highlighted in red. The results demonstrate that Dolphin advances the state-of-the-art in neurosymbolic learning by scaling to more complex problems, e.g., larger versions of HWF, Path, and Mugen, that are beyond the reach of the other baselines which time out within 10 hours. Even for the other benchmarks where baselines do not time out, Dolphin achieves a scalability factor of up to 47x, 62x, and 3.49x against Scallop, ISED, and IndeCateR+ with an average speed up of 13.95x across all baselines for all benchmarks.
We also use the training times per epoch to calculate a scalability factor (Table 5, Appendix D). We see that among cases where baselines timeout, is up to 280x for HWF-19 and is 40.6x faster on average. This results in Dolphin effectively training for more epochs in less time compared with the other baselines, which also allows it to converge to higher accuracies, as we see in Figure 3. We expand on these results in the next RQ.
@@ 3.4 RQ2: Accuracy
Figure 3 presents the accuracy of Dolphin and the baselines on the different benchmarks trained for up to 10 hours. Dolphin accuracies are marked in blue. In all cases, for Dolphin, we report the accuracies of the best-performing provenance. We use the DAMP provenance for MNIST, CLUTRR, and Mugen benchmarks, and the DTKP-AM provenance for the HWF and PathFinder benchmarks.
We observe that in all cases, Dolphin achieves state-of-the-art accuracy among general-purpose neurosymbolic frameworks, except in CLUTRR, where Dolphin’s accuracy is slightly lower than Scallop’s. As we scale up to larger versions of the benchmarks, Dolphin achieves better accuracy, because the baselines either report lower accuracy due to the complexity of the benchmark (e.g., black-box sampling techniques such as ISED on HWF) or fail to converge within 10 hours due to slower per-epoch train time (e.g., Scallop on PathFinder-256, IndeCateR+ on HWF-19). As a result, Dolphin significantly outperforms the second-best benchmark on the largest dataset versions, achieving up to a 20% gain on HWF and 33% on Path. In some cases, given no timeout, Scallop and IndeCateR+ eventually converge to accuracies comparable to Dolphin, as we show in Appendix D.2. However, doing so requires significantly more training time, as discussed earlier in Section 3.3.
These results show that not only do Dolphin’s scalability improvements not come at the cost of accuracy, but Dolphin enables SOTA accuracy when previously unattainable.
@@ 3.5 RQ3: Provenance Comparisons
4 Discussion
Dolphin can be used for any task where the output of a model can be cast as a distribution over probabilities. This abstraction naturally encompasses a wide range of discriminative models in machine learning, such as classifiers, structured prediction systems, and vision-language models. By associating each symbolic object with a probability distribution over its possible values (tags), Dolphin enables downstream symbolic reasoning over uncertain predictions made by neural networks.
For instance, consider an autonomous driving scenario where a standard object detector such as Faster R-CNN outputs bounding boxes and class probabilities. One can represent each detection as a symbolic object whose attributes include coordinates and a Distribution over class probabilities:
Now, consider checking whether a detected person is inside a detected car using symbolic reasoning:
This computation yields a Distribution over True and False, representing the likelihood that two detected objects constitute a person inside a car. This paradigm can generalize to many domains where neural predictions must be interpreted symbolically—for example, relationship inference in vision scenes, symbolic post-processing over structured outputs, or probabilistic parsing in NLP. Dolphin thus offers a principled and composable way to integrate symbolic logic over deep model outputs, opening the door for broader real-world applicability.
@@ 5 Related Work
Scaling techniques. Various methods exist to scale differentiable reasoning. LYRICS (Marra et al., 2019), Logic Tensor Networks (Badreddine et al., 2022), and Tensorlog (Cohen et al., 2020) compile first-order logic constraints into GPU-compatible computation graphs. Other techniques, such as Greedy NTP (Minervini et al., 2020a) and the conditional theorem prover (Minervini et al., 2020b), optimize proof search using heuristics. SLASH (Skryagin et al., 2024) integrates neural networks and probabilistic circuits with Answer Set Programming, achieving scalability by dynamically pruning stochastically insignificant parts of programs. A-NESI (van Krieken et al., 2023) uses learned neural models to approximate the exact probabilistic semantics of WMC, boosting scalability. However, these methods are often task-specific and lack generalizability to broader neurosymbolic learning, or they struggle to scale due to memory constraints when grounding symbolic computations on GPU.
Specialized neurosymbolic solutions. There are many specialized solutions for various neurosymbolic tasks. For instance, NGS (Li et al., 2020) uses a hand-coded syntax to specify the structure of mathematical expressions for HWF. More general solutions, such as NS-CL (Mao et al., 2019) includes a framework for visual question answering that learns symbolic representations for text and images. NeRd (Chen et al., 2021) transforms questions in natural language into executable programs based on symbolic information extracted from text. (Orvieto et al., 2023) proposes a recurrent neural network architecture that achieves 95% accuracy on Path-32 and 94% on Path-128. In contrast, Dolphin is a general system that tries to scale diverse neurosymbolic programs.
@@ 6 Conclusion and Limitations
We proposed Dolphin, a framework for scaling neurosymbolic learning. Dolphin provides abstractions for writing symbolic programs along with pluggable vectorized provenances to compute symbolic gradients. This allows users to write differentiable symbolic programs in Python within PyTorch pipelines that can scale to complex programs and large datasets. We show that Dolphin scales significantly better than existing neurosymbolic frameworks while achieving state-of-the-art performance on a variety of tasks.
@@
Acknowledgements
We thank the reviewers for their insightful feedback that helped to improve this paper. This research was supported by the ARPA-H program on Safe and Explainable AI under the award D24AC00253-00, the NSF award #2313010, and a Google PhD Fellowship.
@@ A.1 WMC approximation
Under both DTKP-WMC and DTKP-AM, the probability of is:
@@ D.1 Comparison of Per Epoch Training Times
D.2 Accuracy without Timeouts
While Figure 3 shows the accuracy of baselines run until a timeout of 10 hours, Figure 5 shows their accuracy when run until convergence or until the test-time accuracy plateaus. We see that only IndeCateR+ for HWF-19 and Scallop for Path-256 are able to match Dolphin’s accuracies. Scallop and ISED are unable to converge for the larger versions of HWF, while ISED is unable to converge for the larger versions of SumN.
D.3 Effect of on Accuracy and Runtime
We report preliminary results on the effect of the top- value in DTKP-AM across different tasks. As shown in Tables 6 and 7, increasing generally has little effect on final accuracy or per-epoch training time. HWF-19 is the only task where increasing offers noticeable gains. For all other benchmarks, accuracy remains stable and runtime scales sub-linearly due to DTKP-AM’s vectorized implementation.
| Benchmark | ||||
| Sum-15 | 9.61 | 10.81 | 10.51 | 10.21 |
| HWF-19 | 8.94 | 99.15 | 96.89 | 95.75 |
| Path-256 | 81.39 | 82.14 | 80.86 | 82.38 |
| CLUTRR-4 | 53.62 | 48.52 | 50.35 | 48.17 |
| Mugen-5K | (94.1 / 95.7) | (95.4 / 95.7) | (95.3 / 95.4) | (95.4 / 95.4) |
| Benchmark | ||||
| Sum-15 | 37.21 | 47.70 | 53.52 | 58.54 |
| HWF-19 | 1.21e3 | 1.40e3 | 1.33e3 | 1.46e3 |
| Path-256 | 1.97e3 | 2.34e3 | 2.10e3 | 2.12e3 |
| CLUTRR-4 | 240.50 | 257.89 | 261.31 | 257.99 |
| Mugen-5K | 460.38 | 464.68 | 470.05 | 465.10 |
D.4 MNIST Sum-N
For this task, the base neural network model is a standard CNN (a 3-layer convolutional network with ReLU activations) classifying each image into 10 classes of digits (0, 1, , 9). The symbolic module sums the Distribution objects over the logits output by the neural model for each image.
@@ D.4 MNIST Sum-N
D.5 Hand-Written Formula